OpenStudy (anonymous):

Simplify the following expression; express in terms of i. square root of -72

OpenStudy (anonymous):

72i

OpenStudy (anonymous):

since \[i =\sqrt{-1}\]

OpenStudy (anonymous):

what why is that the answer

OpenStudy (anonymous):

express in terms of i..... \[i=\sqrt{-1}\] \[72i=\sqrt{-72}\]

OpenStudy (anonymous):

are you sure because it is always a factor of 72

OpenStudy (anonymous):

ohhh.. i'm sorry. i didn't think....

OpenStudy (anonymous):

\[\sqrt{-72}=\sqrt{-1*2*36}=6i \sqrt{2}\]

OpenStudy (anonymous):

this is the right answer.

OpenStudy (anonymous):

ok thank you i thought it looked wrong to me